What is the area of a triangle that has a base of 5 m and a height of 8 m?
Found 2 solutions by jim_thompson5910, moriton:
Answer by jim_thompson5910(35256)

Area = (base*height)/2

A = (bh)/2

A = (5*8)/2

A = 40/2

A = 20


So the area of the triangle is 20 square meters.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by moriton(2)

bxh/2=5*8/2=20 m